OpenStudy (anonymous):

x + 3 / x +5 + x / 6 = 5 / 6

OpenStudy (anonymous):

\[\frac{ x + 3 }{ x + 5 } + \frac{ x }{ 6 } = \frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

find some common denominators so you can add them together

OpenStudy (anonymous):

6(x + 5)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What next?

OpenStudy (anonymous):

\[\frac{ 6(x+3) }{ 6(x+5) }+\frac{ x(x+5) }{ 6(x+5) } = \frac{ 5(x+5) }{ 6(x+5) }\]

OpenStudy (anonymous):

Then you cancel right?

OpenStudy (anonymous):

so now you can just get rid of all the denominators, so: 6(x+3) + x(x+5) = 5(x+5)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

so now just solve for x like a normal quadratic equation

OpenStudy (anonymous):

Hold on one sec.

OpenStudy (anonymous):

6x + 18 + x^2 + 5x = 5x + 25

OpenStudy (anonymous):

Then,

OpenStudy (anonymous):

then move it all on to one side, and solve the quadratic equation

OpenStudy (anonymous):

x^2 + x - 7 = 0

OpenStudy (anonymous):

Is that one correct?

OpenStudy (anonymous):

x^2 +11x +18 = 5x + 25 which equals x^2 + 6x - 7 = 0

OpenStudy (anonymous):

you forgot there was an extra '5x' on the left side

OpenStudy (anonymous):

oh, So then you factor the quadratic equation. hold on.

OpenStudy (anonymous):

(x + 7)(x - 1) = 0

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

looks right to me so x = -7 and 1

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

you're welcome